# Instance-Optimal Imprecise Convex Hull

### v3

@@ 5.2 Speeding up our reconstruction algorithm
The proof is identical to that of Lemma 4.3, with one change: Instead of traversing $\mathbb{E}(\nu)$ for $O(k)$ steps, we do a lookup in $\Gamma(\nu)$ to find the compressed candidate chain that contains $(s,t)$. Let $e$ and $e^{\prime}$ be the first and last candidate chain in this object. Then, all edges in between $e$ and $e^{\prime}$ are candidate edges and hence dividing, so we only need to find the predecessor of $e$ and successor of $e^{\prime}$ in $\mathbb{E}(\nu)$, and check that these edges each are dividing and have endpoints in two different regions.
@@
By Observation 4.1, we obtain two unique areas $R(s),R(t)$ with $s\in V(R(s))$ and $t\in V(R(t))$ in constant time. The edge $(s,t)$ is occupied if and only if $\operatorname{band}(R(s),R(t))$ contains a vertex in $V(F-R(s)-R(t))-\{s\}-\{t\}$. Given $H(R(s))$ and $H(R(t))$, we obtain the edges of $\operatorname{band}(R(s),R(t))$ as a balanced binary tree in cyclical ordering in $O(\log k)$ time through the bridge-finding algorithm [20]. Since $R(s)$ and $R(t)$ are vertically separated, we may split $\operatorname{band}(R(s),R(t))$ into two convex areas: $C(s)$ containing $H(R(s))$ and $C(t)$ containing $H(R(s))$. The edge $(s,t)$ is occupied if and only if one of two conditions holds:
@@
If $R(s)$ is not a point region then $s$ is uniquely a vertex of $R(s)$. What remains is to test whether $C(s)$ contains a vertex in $V(F-R(s))$. Let $\sigma$ be the bit-string corresponding to the ID of $R(s)$ and let $\sigma[i]$ denote inverse of the $i$’th bit in $\sigma$. Note that for all $i\in[\log n]$, $R(s)\not\in F(i,\sigma[i])$ and that for all $R_{b}\neq R(s)$ there exists an integer $i$ such that $R_{b}\in F(i,\sigma[i])$.
@@
The proof is identical to Lemma 4.13, using Lemmas 5.6 and 5.8.
@@
The proof is identical to Corollary 4.15, using Lemma 5.10.
@@
We invoke Lemma 5.6 to find the at most two contiguous subchains $C$ of $\mathbb{E}(\nu)$ with $s,t\in C$, Given any candidate chain $C=(q,\ldots,s,\ldots,r)$, we show how to test in $O(\log m)$ time whether it is spanning in $F$. We apply Observation 4.1 to obtain three unique areas $R(q)$, $R(s)$ and $R(r)$. The chain is spanning if and only if $R(s)$ intersects the inside of $\operatorname{OCH}(R(q),R(r))$. Since $R(q)$ and $R(r)$ are vertically separated and we store their convex hulls in a data structure, we may obtain $\operatorname{OCH}(R(q)\cup R(r))$ as a balanced binary tree in $O(\log m)$ time. We then do convex hull intersection testing between the convex hull of $R(s)$ and $\operatorname{OCH}(R(q)\cup R(r))$ in $O(\log m)$ time.
@@
This proof is identical to that of Corollary 4.11, invoking Lemma 5.14 instead of Lemma 4.9.
@@ 6 Pairwise disjoint disks with a radius in $[1,k]$, or ply at most $k$
[…]
@@
[…]
@@
Corollary 6.5 leads to the following data structure.
@@
We dynamically maintain $\operatorname{MCD}(F)$ subject to retrievals in $F$. We then show that we can augment the Median Cut Decomposition so that each node $\nu\in\operatorname{MCD}(F)$ also maintains a concatenable queue of $\operatorname{OCH}(F(\nu))$. Finally, we use this data structure to obtain an instance-optimal reconstruction strategy in a manner analogous to Section 4.
@@ 6.1 Dynamically maintaining $\operatorname{MCD}(F)$
We want to maintain the property that $\operatorname{MCD}(F)$ is a balanced binary tree. Many schemes for balancing binary trees rely on tree rotations. However, when rotating around a node $\nu$ (making it the parent node of its parent) the family of regions $F(\nu)$ (and thereby $\phi(\nu)$) may have $O(n)$ recourse. To avoid this issue, we use scapegoating, introduced by Andersson [2]. Instead of using rotations to maintain height-balance, scapegoating rebuilds the balanced tree on $\nu$ from scratch whenever the height of a node $\nu$ becomes too large compared to the size of $\nu$. If constructing the tree takes $O(n\cdot f(n))$ time, then insertions under scapegoating have $O(f(n)\log n)$ amortized update cost. We briefly note that when constructing $\operatorname{MCD}(F)$ as a perfectly balanced tree, the amortised potential in each node is zero. This implies:
@@ 6.2 Augmenting $\operatorname{MCD}(F)$
We consider $C_{x},C_{y}$ and $C_{\nu}^{*}$ from Definition 6.10. By Lemma 6.1 (or, Lemma 6.3), $C_{x},C_{y}$ and $C_{\nu}$ have $O(k)$ edges, which we obtain in $O(k+\log n)$ time. We store a copy of $C_{x}$ and $C_{y}$ as a balanced binary tree. We then apply merge sort and get $C^{*}=\operatorname{OCH}(C_{x}\cup C_{y}\cup C_{\nu}^{*})$ as a balanced binary tree in $O(k)$ time.
@@
Lemma 6.11 implies an $O(n\cdot(k+\log n))$ time construction algorithm for our augmented $\operatorname{MCD}(F)$. Thus, after inserting $\lambda$, we may use the scapegoating amortised rebuilding technique to maintain the balance of $\operatorname{MCD}(F)$ in $O(k\log n+\log^{2}n)$ amortised time. Suppose our amortized rebuilding scheme rebuilds some subtree $\mu^{*}$ that contains $\lambda$ (if no subtrees were rebuilt, $\mu^{*}$ is the parent of $\lambda$). Let $\pi$ denote the path in $\operatorname{MCD}(F)$ from the root to $\mu^{*}$. What remains is to restore $\operatorname{MCD}(F)$ along $\pi$.
We invoke Lemma 6.13 over the top-down traversal of $\pi$. This way, we obtain for each $\nu\in\pi$, with children $(\mu,\gamma)$, the trees $\mathbb{E}(\mu)$ and $\mathbb{E}(\gamma)$ before the update. If $\lambda$ is to be in the subtree rooted at $\mu$, we keep $\mathbb{E}(\gamma)$ in memory and invoke Lemma 6.11 on $\mu$ next. We do this until we arrive at $\mu^{*}$, which has one new child and one unchanged child node, and we start a bottom-up traversal. Let $\nu\in\pi$ be the next node in our bottom-up traversal and let $\nu$ have children $x$ and $y$. Given the up-to-date trees $\mathbb{E}(x)$, $\mathbb{E}(y)$ and $\operatorname{CH}(\phi(\nu))$ we invoke Lemma 6.11 to compute $(\mathbb{E}(\nu),\mathbb{E}^{*}(x),\mathbb{E}^{*}(y))$ in $O(k+\log n)$ time. Since $\pi$ as length at most $O(\log n)$, the remaining update takes $O(k\log n+\log^{2}n)$ time.
@@
From here-on, we use our data structure in an identical manner to Section 4. Let $\nu\in\operatorname{MCD}(F)$. We augment the concatenable queues of a node $\nu$ by storing four additional balanced trees:
@@
For an update at a node $\mu$, only the $O(\log n)$ nodes on the root-to-leaf path of $\mu$ can gain or lose a contiguous subchain of $\mathbb{E}(\nu)$ that is spanning in $F$. For each such node $\nu$ with children $x,y$, we invoke Lemma 6.17 to find all $O(k)$ candidate chains that are contiguous subchains of $\operatorname{CH}(\nu)$ but not of $\operatorname{CH}(x)$ or $\operatorname{CH}(y)$. We use Lemma 6.21 to check which of these chains are spanning in $F$. We then maintain the balanced binary tree associated to each concatenable queue $\mathbb{E}^{*}(\nu)$ in an identical manner as previous corollaries. This takes $O(k\log n)$ time per update. During scapegoating, we consider all $O(n)$ nodes $\nu$, so this takes amortized $O(k\log n)$ time too.
@@
Any vertex of $\operatorname{OCH}(F-R_{a}-R_{b})$ must also be a vertex that appears in a chain in $\bigcup\{L_{a},L_{b},\Psi_{a},\Psi_{b},H_{a},H_{b}\}$. If follows that $(s,t)$ is occupied if and only if the convex hull of $(R_{a},R_{b})$ contains a vertex in a chain in $\bigcup\{L_{a},L_{b},\Psi_{a},\Psi_{b},H_{a},H_{b}\}$. There are $O(\log n)$ convex chains in this set. For each of these chains we use the standard algorithm to test whether two convex hulls intersect in $O(\log n)$ time [7].
@@
We invoke Lemma 6.17 to find all $O(k)$ candidate chains that contain a bridge of $\nu$. We extend each such chain by one edge to the left or right to a chain $C^{\prime}$, and use Lemma 6.25 to check whether $C^{\prime}$ is hit in $F$.
@@
For an update at a node $\mu$, only the $O(\log n)$ nodes on the root-to-leaf path of $\mu$ can gain or lose a contiguous subchain of $\mathbb{E}(\nu)$ that is hit in $F$. For each such node $\nu$ with children $x,y$, we invoke Lemma 6.27 to find all $O(k)$ chains that are hit in $F$ and that are contiguous subchains of $\operatorname{CH}(\nu)$ but not of $\operatorname{CH}(x)$ or $\operatorname{CH}(y)$. We then maintain the balanced binary tree associated to each concatenable queue $\mathbb{E}^{*}(\nu)$ in an identical manner as previous corollaries. This takes $O(k\log^{3}n)$ time per update. During scapegoating, we consider all $O(n)$ nodes $\nu$, so this takes amortized $O(k\log^{3}n)$ time too.
Through Corollary 6.19, 6.29, and 6.23 we always have at the root of $\operatorname{MCD}(F)$ a doubly linked list of all edges of $\operatorname{OCH}(F)$ that are non-dividing, and a doubly linked list of all continuous subchains of $\operatorname{OCH}(F)$ that are spanning in $F$, or hit in $F$. We observe that Algorithm 1 only considers occupied edges if all edges on $\operatorname{OCH}(F)$ are dividing, in which case any occupied edge corresponds to a hit subchain. This implies that we may execute Algorithm 1 and so:
@@
Appendix A Further related work
We elaborate on the two related works most closely aligned with this paper. Bruce, Hoffmann, Krizanc, and Raman [3] presented the first instance-optimal retrieval strategy. However, they did not provide an explicit retrieval program with an accompanying algorithmic running time. We elaborate on their retrieval strategy and discuss the complexity considerations that arise when implementing a corresponding retrieval program.
Recently and independently of our work, Löffler and Raichel [18] developed an algorithm for reconstructing the convex hull when $F$ is a set of unit disks. They restrict the overlap in $F$ to ply $k$. The number of retrievals performed by their algorithm is not instance-optimal. We discuss their approach and provide commentary on its running time.
The algorithm by Bruce, Hoffmann, Krizanc, and Raman [3] and its running time.
Bruce et al. [3] present an instance-optimal retrieval strategy based on categorising regions in $F$. The key idea is to partition the curent set of regions into four categories (see Figure 9). Each region is assigned the highest applicable category, as follows:

### v4

@@ 5.2 Speeding up our reconstruction algorithm
The proof is identical to that of Lemma˜4.3, with one change: Instead of traversing $\mathbb{E}(\nu)$ for $O(k)$ steps, we do a lookup in $\Gamma(\nu)$ to find the compressed candidate chain that contains $(s,t)$. Let $e$ and $e^{\prime}$ be the first and last candidate chain in this object. Then, all edges in between $e$ and $e^{\prime}$ are candidate edges and hence dividing, so we only need to find the predecessor of $e$ and successor of $e^{\prime}$ in $\mathbb{E}(\nu)$, and check that these edges each are dividing and have endpoints in two different regions.
@@
By Observation 4.1, we obtain two unique areas $R(s),R(t)$ with $s\in V(R(s))$ and $t\in V(R(t))$ in constant time. The edge $(s,t)$ is occupied if and only if $\operatorname{band}(R(s),R(t))$ contains a vertex in $V(F-R(s)-R(t))-\{s\}-\{t\}$. Given $H(R(s))$ and $H(R(t))$, we obtain the edges of $\operatorname{band}(R(s),R(t))$ as a balanced binary tree in cyclical ordering in $O(\log k)$ time through the bridge-finding algorithm [DBLP:conf/stoc/OvermarsV80]. Since $R(s)$ and $R(t)$ are vertically separated, we may split $\operatorname{band}(R(s),R(t))$ into two convex areas: $C(s)$ containing $H(R(s))$ and $C(t)$ containing $H(R(s))$. The edge $(s,t)$ is occupied if and only if one of two conditions holds:
@@
If $R(s)$ is not a point region then $s$ is uniquely a vertex of $R(s)$. What remains is to test whether $C(s)$ contains a vertex in V(FR(s))V(F-R(s)). Let $\sigma$ be the bit-string corresponding to the ID of $R(s)$ and let $\sigma[i]$ denote inverse of the $i$’th bit in $\sigma$. Note that for all $i\in[\log n]$, $R(s)\not\in F(i,\sigma[i])$ and that for all $R_{b}\neq R(s)$ there exists an integer $i$ such that $R_{b}\in F(i,\sigma[i])$.
@@
The proof is identical to Lemma˜4.13, using Lemmas˜5.6 and 5.8.
@@
The proof is identical to Corollary˜4.15, using Lemma˜5.10.
@@
We invoke Lemma˜5.6 to find the at most two contiguous subchains $C$ of $\mathbb{E}(\nu)$ with $s,t\in C$, Given any candidate chain $C=(q,\ldots,s,\ldots,r)$, we show how to test in $O(\log m)$ time whether it is spanning in $F$. We apply Observation 4.1 to obtain three unique areas $R(q)$, $R(s)$ and $R(r)$. The chain is spanning if and only if $R(s)$ intersects the inside of $\operatorname{OCH}(R(q),R(r))$. Since $R(q)$ and $R(r)$ are vertically separated and we store their convex hulls in a data structure, we may obtain $\operatorname{OCH}(R(q)\cup R(r))$ as a balanced binary tree in $O(\log m)$ time. We then do convex hull intersection testing between the convex hull of $R(s)$ and $\operatorname{OCH}(R(q)\cup R(r))$ in $O(\log m)$ time.
@@
This proof is identical to that of Corollary˜4.11, invoking Lemma˜5.14 instead of Lemma˜4.9.
@@ 6 Pairwise disjoint disks with a radius in $[1,k]$, or ply at most $k$
For $F$ a family of pairwise disjoint disks with radii in $[1,k]$, $A\subseteq F$, and $S$ a vertical slab of width $2k$, let $F^{\prime}=\{R\in F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptA\mid R\subset S\}$. Then $\operatorname{OCH}(F^{\prime})$ has $O(k)$ edges.
@@
For $F$ a family of unit-size disks of ply $k$, let $A\subseteq F$, and $S$ be a vertical slab of width $2$, let $F^{\prime}=\{R\in F\hskip-1.42271pt\odot_{\hskip-1.42271pt\textit{P}}\hskip-0.85355ptA\mid R\subset S\}$. Then $\operatorname{OCH}(F^{\prime})$ has $O(k)$ edges.
@@
Corollary˜6.5 leads to the following data structure.
@@
We dynamically maintain $\operatorname{MCD}(F)$ subject to retrievals in $F$. We then show that we can augment the Median Cut Decomposition so that each node $\nu\in\operatorname{MCD}(F)$ also maintains a concatenable queue of $\operatorname{OCH}(F(\nu))$. Finally, we use this data structure to obtain an instance-optimal reconstruction strategy in a manner analogous to Section˜4.
@@ 6.1 Dynamically maintaining $\operatorname{MCD}(F)$
We want to maintain the property that $\operatorname{MCD}(F)$ is a balanced binary tree. Many schemes for balancing binary trees rely on tree rotations. However, when rotating around a node $\nu$ (making it the parent node of its parent) the family of regions $F(\nu)$ (and thereby $\phi(\nu)$) may have $O(n)$ recourse. To avoid this issue, we use scapegoating, introduced by Andersson [andersson1989improving]. Instead of using rotations to maintain height-balance, scapegoating rebuilds the balanced tree on $\nu$ from scratch whenever the height of a node $\nu$ becomes too large compared to the size of $\nu$. If constructing the tree takes $O(n\cdot f(n))$ time, then insertions under scapegoating have $O(f(n)\log n)$ amortized update cost. We briefly note that when constructing $\operatorname{MCD}(F)$ as a perfectly balanced tree, the amortised potential in each node is zero. This implies:
@@ 6.2 Augmenting $\operatorname{MCD}(F)$
We consider $C_{x},C_{y}$ and $C_{\nu}^{*}$ from Definition 6.10. By Lemma˜6.1 (or, Lemma 6.3), $C_{x},C_{y}$ and $C_{\nu}$ have $O(k)$ edges, which we obtain in $O(k+\log n)$ time. We store a copy of $C_{x}$ and $C_{y}$ as a balanced binary tree. We then apply merge sort and get $C^{*}=\operatorname{OCH}(C_{x}\cup C_{y}\cup C_{\nu}^{*})$ as a balanced binary tree in $O(k)$ time.
@@
Lemma˜6.11 implies an $O(n\cdot(k+\log n))$ time construction algorithm for our augmented $\operatorname{MCD}(F)$. Thus, after inserting $\lambda$, we may use the scapegoating amortised rebuilding technique to maintain the balance of $\operatorname{MCD}(F)$ in $O(k\log n+\log^{2}n)$ amortised time. Suppose our amortized rebuilding scheme rebuilds some subtree $\mu^{*}$ that contains $\lambda$ (if no subtrees were rebuilt, $\mu^{*}$ is the parent of $\lambda$). Let $\pi$ denote the path in $\operatorname{MCD}(F)$ from the root to $\mu^{*}$. What remains is to restore $\operatorname{MCD}(F)$ along $\pi$.
We invoke Lemma˜6.13 over the top-down traversal of $\pi$. This way, we obtain for each $\nu\in\pi$, with children $(\mu,\gamma)$, the trees $\mathbb{E}(\mu)$ and $\mathbb{E}(\gamma)$ before the update. If $\lambda$ is to be in the subtree rooted at $\mu$, we keep $\mathbb{E}(\gamma)$ in memory and invoke Lemma˜6.11 on $\mu$ next. We do this until we arrive at $\mu^{*}$, which has one new child and one unchanged child node, and we start a bottom-up traversal. Let $\nu\in\pi$ be the next node in our bottom-up traversal and let $\nu$ have children $x$ and $y$. Given the up-to-date trees $\mathbb{E}(x)$, $\mathbb{E}(y)$ and $\operatorname{CH}(\phi(\nu))$ we invoke Lemma˜6.11 to compute $(\mathbb{E}(\nu),\mathbb{E}^{*}(x),\mathbb{E}^{*}(y))$ in $O(k+\log n)$ time. Since $\pi$ as length at most $O(\log n)$, the remaining update takes $O(k\log n+\log^{2}n)$ time.
@@
From here-on, we use our data structure in an identical manner to Section˜4. Let $\nu\in\operatorname{MCD}(F)$. We augment the concatenable queues of a node $\nu$ by storing four additional balanced trees:
@@
For an update at a node $\mu$, only the $O(\log n)$ nodes on the root-to-leaf path of $\mu$ can gain or lose a contiguous subchain of $\mathbb{E}(\nu)$ that is spanning in $F$. For each such node $\nu$ with children $x,y$, we invoke Lemma˜6.17 to find all $O(k)$ candidate chains that are contiguous subchains of $\operatorname{CH}(\nu)$ but not of $\operatorname{CH}(x)$ or $\operatorname{CH}(y)$. We use Lemma˜6.21 to check which of these chains are spanning in $F$. We then maintain the balanced binary tree associated to each concatenable queue $\mathbb{E}^{*}(\nu)$ in an identical manner as previous corollaries. This takes $O(k\log n)$ time per update. During scapegoating, we consider all $O(n)$ nodes $\nu$, so this takes amortized $O(k\log n)$ time too.
@@
Any vertex of $\operatorname{OCH}(F-R_{a}-R_{b})$ must also be a vertex that appears in a chain in $\bigcup\{L_{a},L_{b},\Psi_{a},\Psi_{b},H_{a},H_{b}\}$. If follows that $(s,t)$ is occupied if and only if the convex hull of $(R_{a},R_{b})$ contains a vertex in a chain in $\bigcup\{L_{a},L_{b},\Psi_{a},\Psi_{b},H_{a},H_{b}\}$. There are $O(\log n)$ convex chains in this set. For each of these chains we use the standard algorithm to test whether two convex hulls intersect in $O(\log n)$ time [chazelle1980detection].
@@
We invoke Lemma˜6.17 to find all $O(k)$ candidate chains that contain a bridge of $\nu$. We extend each such chain by one edge to the left or right to a chain $C^{\prime}$, and use Lemma˜6.25 to check whether $C^{\prime}$ is hit in $F$.
@@
For an update at a node $\mu$, only the $O(\log n)$ nodes on the root-to-leaf path of $\mu$ can gain or lose a contiguous subchain of $\mathbb{E}(\nu)$ that is hit in $F$. For each such node $\nu$ with children $x,y$, we invoke Lemma˜6.27 to find all $O(k)$ chains that are hit in $F$ and that are contiguous subchains of $\operatorname{CH}(\nu)$ but not of $\operatorname{CH}(x)$ or $\operatorname{CH}(y)$. We then maintain the balanced binary tree associated to each concatenable queue $\mathbb{E}^{*}(\nu)$ in an identical manner as previous corollaries. This takes $O(k\log^{3}n)$ time per update. During scapegoating, we consider all $O(n)$ nodes $\nu$, so this takes amortized $O(k\log^{3}n)$ time too.
Through Corollary˜6.19, 6.29, and 6.23 we always have at the root of $\operatorname{MCD}(F)$ a doubly linked list of all edges of $\operatorname{OCH}(F)$ that are non-dividing, and a doubly linked list of all continuous subchains of $\operatorname{OCH}(F)$ that are spanning in $F$, or hit in $F$. We observe that Algorithm 1 only considers occupied edges if all edges on $\operatorname{OCH}(F)$ are dividing, in which case any occupied edge corresponds to a hit subchain. This implies that we may execute Algorithm 1 and so:
@@
Appendix A Related work
We elaborate on the two related works most closely aligned with this paper. Bruce, Hoffmann, Krizanc, and Raman [bruce2005efficient] presented the first instance-optimal retrieval strategy. However, they did not provide an explicit retrieval program with an accompanying algorithmic running time. We elaborate on their retrieval strategy and discuss the complexity considerations that arise when implementing a corresponding retrieval program.
Recently and independently of our work, Löffler and Raichel [loffler2025preprocessing] developed an algorithm for reconstructing the convex hull when $F$ is a set of unit disks. They restrict the overlap in $F$ to ply $k$. The number of retrievals performed by their algorithm is not instance-optimal. We discuss their approach and provide commentary on its running time.
The algorithm by Bruce, Hoffmann, Krizanc, and Raman [bruce2005efficient] and its running time.
Bruce et al. [bruce2005efficient] present an instance-optimal retrieval strategy based on categorising regions in $F$. The key idea is to partition the curent set of regions into four categories (see Figure 9). Each region is assigned the highest applicable category, as follows: